OpenStudy (thecalchater):

Locate the absolute value extrema of the function f(x)=cos(pi x) on the closed interval [0,3/4] Determine whether Rolle's Theorem can be applied to the function f(x)=(x+3)(x-2)^2 on the closed interval [-3,2] if rolle's Theorem can be applied find all numbers c in the open interval (-3.2) such that f' (c)=0 Determine whether the Mean Value Theorem can be applied to the function f(x)=2 sin x + sin 2x on the closed interval [7 pi, 8 pi] If the mean value theorem can be applied find all numbers c in the open interval (7 pi, 8 pi) such that f'(c)= f(8 pi)-f(7 pi)/8 pi-7 pi

OpenStudy (thecalchater):

@satellite73

OpenStudy (thecalchater):

@ganeshie8

OpenStudy (thecalchater):

I figured our rolles theorem but not the MVT i got the derivative and then plugged in 7pi and 8pi and got 0 where do i go from there?

OpenStudy (thecalchater):

@ganeshie8 @TheSmartOne

OpenStudy (anonymous):

For the first, you can compute the derivative, find the critical points, and apply the derivative test: \[f(x)=\cos\pi x\implies f'(x)=-\pi\sin\pi x=0~~\implies \pi x=n\pi\implies x=n\]where \(n\) is any integer. How many integers exist in the interval \(\left[0,\dfrac{3}{4}\right]\)? For the second, \(f(x)\) must be continuous and differentiable over the given interval \((a,b)\), with \(f(a)=f(b)\). Since this is true, this guarantees the existence of \(c\) such that \[f'(c)=(c-2)^2+2(c+3)(c-2)=0\implies c=\cdots\] For the third, the same conditions must hold for \(f(x)\), except that \(f(a)=f(b)\). Since \(\sin x\) is a continuous function and any linear combination of continuous functions must also be continuous, you know that \(f(x)\) satisfies the MVT. Now you want to find \(c\) such that \[f'(c)=\frac{f(8\pi)-f(7\pi)}{8\pi-7\pi}\implies 2\cos c+2\cos2c=0\]To solve for \(c\), use an identity: \[\cos2x=\cos^2x-\sin^2x=2\cos^2x-1\]So you have this equation with the LHS quadratic in \(\cos c\): \[2\cos c+2\cos2c=4\cos^2c+2\cos c-2=0\]

OpenStudy (thomas5267):

Why does the second question requires the Rolle's theorem. Since [-3,2] is closed and bounded and the f(x) in second question is continuous, just use the extreme value theorem.